OpenStudy (anonymous):

I'm having trouble finding what c equals for 6-5c=-29

OpenStudy (anonymous):

c = 7

OpenStudy (cwrw238):

subtract 6 from both sides 6-6-5c = -29-6 -5c = -35 now divide both sides by -5 to find c

OpenStudy (anonymous):

6 - 5c = -29 6-6-5c = -29-6 -5c = -35 -5/-5c = -35/-5 c = 7

OpenStudy (anonymous):

Thank you guys so much

OpenStudy (cwrw238):

yw

OpenStudy (anonymous):

np :)